YALO Observatory
Telescope & Instrument News

Updated: 2002 August 16

This page provides news about the YALO Telescope, ANDICAM, and the web pages of interest to observers, proposers, and consortium members. Entries are in reverse chronological order (most recent news first).


2002 Sept 24: YALO Project End of Operations

2002 June 27: CCD Problems - an Update
2002 Feb 8: Back in Operation
2002 Jan 28: ANDICAM Offline: CCD Shutter Dead
2002 Jan 14: CCD Shutter Problems
2002 Jan 14: Web Server Downtime
2001 Dec 6: Johnson U replaces Stromgren u
2001 Feb 19: Stromgren Filter FOV
2000 Nov 27: Changes to Report Archives
2000 Sept 28: CCD POSTCLEAN set correctly
2000 Sept 12: New Web Tools
2000 Sept 8: CCD Software Upgrade
2000 Sept 1: Automatic Observing Log Archiving

2002 September 24: YALO Project End of Operations

After more than four successful and highly productive years of operation, the YALO Project was completed and observing operations ended on 2002 September 24. This move will make way for the start of the SMARTS Project, and the move of the ANDICAM to another telescope as part of that venture early next year. This website and its observing records will be kept online for the foreseeable future as a service to all YALO users.

On behalf of all of us who have worked on the YALO project since its inception, we would like to thank all of the astronomers whose imaginative use of this unique facility have made it an unqualified scientific success. We are among the first such projects to demonstrate the scientific utility of this mode of observing, and to do it on a shoe-string budget. The YALO operations model will be the basis of our future ANDICAM operations with SMARTS. We are not so much ending YALO as moving on to the next logical step along these lines.

In particular, the YALO project would not have been possible were it not for the tireless dedication of our two observers, Juan Espinoza and David Gonzalez. Thanks guys!

On to SMARTS...

2002 June 27: CCD Problems - an Update

Shortly after getting back into operation after two weeks of terrible weather at CTIO, the substitute observer Sergio reported that the ANDICAM CCD had stopped working. The last night of CCD data was June 5. On June 6, CCD images read onto disk only contained what looked like noise+bias signal regardless of integration time or degree of illumination, and with none of the usual artifacts from hot areas or blocked columns. We also saw no cosmic rays. We could occasionally get non-bias signal if we blasted the detector with light, but nothing consistent.

By June 11, a series of software tests conducted remotely by OSU personnel with help from Juan Espinoza determined that it was not a problem in the data-taking software. Bruce Atwood was on travel, but nevertheless got in touch with electronic technicians at CTIO and initiated a series of tests over the phone to try to isolate the problem in the CCD readout electronics. By June 14 it was clear that there was no obvious solution, and they had failed to pinpoint the problem on any of the readout boards. While there are no firm conclusions as yet, it is increasingly likely that the CCD detector itself is malfunctioning.

At this date we are looking at options for verifying the fault and replacing the CCD if necessary. The latter requires a significant amount of effort. In the meantime we are proceeding with IR-only observations. More news will be posted when we learn more.

2002 Feb 8: Back in Operation

Bruce Atwood arrived at CTIO on Feb 5 and replaced the shutter on Feb 6. Shutter tests on the night of Feb 6 and afternoon Feb 7 verified that we have a working shutter. Juan executed an old (pre-breakdown) observing program to verify that everything was back to normal.

Regular science operations resumed on the night of February 8.

Many thanks to Bruce Atwood, who made the rescue run, and to Oscar Saa and the technical support staff at CTIO for all of their invaluable assistance!

2002 Jan 28: ANDICAM Offline: CCD Shutter Dead

The CCD camera shutter finally died a definitive death during testing at CTIO. Oscar Saa and Bruce Atwood were unable to revive the shutter, but did determine it was not an electronics failure. The prognosis is a total mechanical failure. ANDICAM will remain offline while Bruce Atwood prepares to travel to CTIO to replace the shutter. The earliest we could be back online is Thursday, Feb 7.

2002 Jan 14: CCD Shutter Problems

The CCD camera shutter in the ANDICAM has been sticking open intermittently. The problem first appeared in mid-December following the recent filter change, but then went away for the most part. Most CCD images of the last few weeks have been unaffected, but when it does strike the problem is kind of obvious.

The main symptom is that you will see bright stars streaking vertically near the middle of the CCD, usually looking like a line of beads extending above and below the star images. The brighter the star the brighter the streak. The cause is the shutter sticking part-way open during readout. The effect is particulary damaging to flats and bias images. [Example of an affected I-band image].

During routine tests to try to determine the nature of the problem and its duty cycle during the Christmas/New Years's break, the problem seemed to have cured itself. A regular regime of tests was instituted, and late during the week of January 6-12 we resumed taking dome flats with the CCD camera.

The problem resurfaced on the night of January 13/14 with a vengeance. Here are images of a flat-field screen using the entire CCD of the shutter stuck open in two positions: Mostly Open and Mostly Closed. A series of tests has been initiated to try to characterize the problem. We at OSU are also working to get a new shutter prepared and to make up detailed instructions on how to replace it should that be required. Since this involves opening the vacuum vessel, we anticipate about 3 days of downtime will be required for a replacement, and want to be sure it is needed before proceeding.

There is no "fix" for images clobbered by this artifact, but since it tends to mostly affect the boundary between the two CCD halves, it does not (usually) affect the region of the CCD that is shared by the IR channel (where the observers are instructed to center the targets).

2002 Jan 14: Web Server Downtime

The YALO web and ftp servers will be unavailable from noon EST on Tuesday, January 15 until 9am EST on Wednesday January 16. The downtime is needed to move the YALO web pages and associated observer files over to the new department web and ftp server and to turn off the old server.

2001 Dec 6: Johnson U Filter Replaces Stromgren u Filter

The Stromgren u-band filter in the CCD filter wheel (position 2) has been replaced with a filter that reproduces the Johnson U bandpass. This filter completes the UBVRI set, and has substantially higher throughput than the Stromgren u filter. [rwp]

All web forms and device tables at the telescope have been updated to reflect the change.

2001 Feb 19: Stromgren Filter FOV

In response to requests from AURA users, I have prepared a "finder" image for the Stromgren filter set on the CCD channel.

The Stromgren filters are only 25-mm diameter, and thus only partially illuminate a round spot centered on the ANDICAM CCD. The effective FOV is 500-arcseconds (1660 pixels), with some vignetting, and a little off center towards the top of the image as viewed by IRAF unflipped.

The orientation of the FOV on the sky is shown in this screenshot of a star field imaged in the Stromgren y filter (300s total integration). This is as the image would be displayed by IRAF (XVista users would see E/W inverted vertically). The circle shows 500-arcsec diameter, and the noisy left half of the CCD has been masked.

Because of the persistent noise problems with the left half of the ANDICAM CCD, users requiring precision photometry should locate their targets in the rigth half of the CCD, within the 500-arcsec circle. [rwp/osu]

2000 Nov 27: Changes to the Report Archives

Security changes to our web server at OSU required me to modify the nightly report archives to create "index" files for the reports. Our server no longer permit access to "bare directories". Nightly observing reports have been indexed, and the reporting scripts modified to re-create the index files as required. Nightly trouble reports will be indexed like this shortly. [rwp/osu]

2000 Sept 28: CCD POSTCLEAN set correctly

In taking sequences of short (20s) CCD images, we noticed that the first image of each set had many more cosmic ray events than all of the preceeding images. On checking the IC settings, we found that POSTCLEAN, which somewhat counter-intuitively sets the number of pre-exposure erase cycles had been set to zero (0). This means that the CCD was not being erased before each image, and if the first image was preceeded by a long period of deadtime, it would have greater cosmic rays and effective dark time.

POSTCLEAN has now been set to 1, which will have a minor (2-3 second) impact upon the total execution time, but well within our canonical 5-10% estimate envelope, so the execution-time calculator has not been changed. You should notice that the first bias frame of a sequence or first short-frame of a sequence not preceeded by another image should be cleaner.

We do not know why this variable was set to zero, or by whom.

2000 Sept 12: New Web Tools

New web tools are available for YALO investigators on the Phase II Observing Program Pages:

  1. Observation Template (Obs) file editor.

  2. Simple observing log search tool.

These tools should simplify making program modifications and tracking the progress of your observations. If you encounter any problems, please contact Rick Pogge (pogge@astronomy.ohio-state.edu).

2000 Sept 8: CCD Software Upgrade: READ THIS!

We have upgraded the software on the CCD channel data-taking system. This institutes a number of changes:

  1. We can now readout the entire 2048x2048 CCD.
  2. Bias and overscan must now take into account 2-amplifier readouts (details below).
  3. The CCD gain has been reset to its correct nominal values, but...
  4. this affects the readout time, and hence all observation execution times need to be revisited for on-going programs.

We'll treat each of these issues in turn.

Full-Frame Readout
Hooray! An error of 1 bit (!) in the CCD sequencer map was responsible for the garbage on the left-half of the ANDICAM CCD. This was found and fixed and we can now readout both halves of the detector. This restores the full ~10-arcmin field-of-view of ANDICAM.

Bias and Overscan Issues
We readout the CCD with two amplifiers, one for the each of the left and right halves of the detector. The horizontal readout registers on the Lick/Loral-3 CCD have 16 extra pixels for each readout amplifier that are read before the data pixels in a line. These constitute "hardware underscan" pixels. After reading each line of data, we readout an additional 32 pixels of "overscan" at the end of each line. These latter are used to measure the DC bias level during the readout of that line on that amplifier, and constitute the traditional "overscan" columns familiar to most CCD observers. Our data-taking system repacks the data and under/overscan pixels so that the non-image pixels are to the outside (left and right) of the data sections. This can make it look like the images have extra overscan. The diagram below should clarify this.

[Schematic of a raw ANDICAM CCD frame showing the locations of the data, underscan, and overscan regions. (18Kb GIF)]
[Typical Bias Image (161Kb GIF) before subtraction of overscan bias. This image is shown with exaggerated contrast, +/-20 ADU, to emphasize the difference in the mean DC bias levels between the two halves of the device. This difference vanishes after overscan subtraction. The pattern noise on the left half is also visible, as are the bright pixels in the first 10-odd data columns at left.]
With 1x1 binning, the raw image size is 2144x2048 pixels, and the data and bias sections for each half of the device are:
      DATASEC1  [49:1072,1:2048]
      BIASSEC1  [1:32,1:2048]
      DATASEC2  [1073:2097,1:2048]
      BIASSEC2  [2113:2144,1:2048]
      TRIMSEC   [49:2096,1:2048]
The BIASSEC coordinates explicitly avoid the 16 pixels of "hardware underscan" in each line, and only use the 32 pixels of "overscan" to determine the DC bias level. Note that the first 10-odd data columns on the left side are unusually bright, and may be discarded with a TRIMSEC of [60:2096:1:2048].

With 2x2 binning the raw image size is 1072x1024 pixels, and the correspondng data and bias sections are:

      DATASEC1  [25:536,1:1024]
      BIASSEC1  [1:16,1:1024]
      DATASEC2  [537:1048,1:1024]
      BIASSEC2  [1057:1072,1:1024]
      TRIMSEC   [25:1048,1:1024]
With 2x2 binning we now have 8 pixels of "underscan" and 16 pixels of "overscan" for each side. The BIASSEC's above avoid the hardware underscan as noted above. In addition, the first 8 data columns are bad in the 2x2 binned images, and may be discarded by using a TRIMSEC of [32:1048,1:1024].

The DC bias level is slightly different from one half to the other, but this should comes out in the overscan processing. However, there is a mysterious pattern noise (ripple) with a peak-to-peak amplitude of ~3-4ADU that is visible in the left-hand side of the detector in bias (zero) images. This will be harder to diagnose fix, and does not entirely come out in 2-D bias (zero) subtraction.

CCD Gain:
Back in November 1999 when we replaced a crashed disk in the CCD controller, we did not notice that an outdated config file was on that disk that inadvertently changed the CCD gain setting. As a consequence, the CCD gain has been running high, ~8 electrons/ADU, since November 1999, resulting in a higher effective readout noise (nearly 20 electrons) due to digitization error. This was discovered in Feb 2000 by Nick Suntzeff, and called to our attention.

We have reset the gain with the software upgrade. The latest gain and readout noise numbers for the ANDICAM CCD are:

gain = 3.6 electrons/ADU
readout noise = 11 electrons (rms)
Please use these for all exposure time estimates.

CCD Readout Times
Changing the gain has the effect of making the readout time longer than before. The new nominal readout overheads are:
1x1 binning: 80 seconds (74s readout + 6s pre-exposure erase/setup)
2x2 binning: 24 seconds (24s readout + 6s pre-exposure erase/setup)
The observing template creation scripts have been updated to reflect these changes.

Observers who have been trying to reduce deadtime in dual-mode (CCD+IR) observations may wish to revise their observing templates as necessary. Note, however, that there is little point in trying to make the deadtime smaller than ~20 seconds per template because of various unpredictable system latencies that can erase any apparent reduction of deadtime.

You can do this most simply by using the new obs file editor. After logging into the Phase II Preparation tools, select "Edit Old Observation Template Files" from the tools menu below, and follow the instructions.

Note that the IR array readout time is unaffected.

2000 Sept 1: Automatic Observing Log Archiving

Starting Sept 1, we are automatically archiving all observing logs on the web at 10am the following morning. These are available in the Observation Log Archive index. Some logs may not be available right away if the network is acting up. Missing logs usually appear the following day (unless one of us intervenes by hand, unlikely after classes start at OSU).

Return to the YALO Observing Page

Updated: 2002 August 16 [rwp/osu]